A 22 g ball of clay traveling east at 3.8 m/s collides with a 38 g ball of clay traveling north at 2.2 m/s .

Part A: What is the movement direction of the resulting 60 g blob of clay? Express your answer in degrees north of east.

Part B: What is the speed of the resulting 60 g blob of clay? Express your answer with the appropriate units.
Answer #1

by conservation of momentum

initial horizontal momentum = final horizontal momentum

22 * 10^-3 * 3.8 = 60* 10^-3 * v * cos(theta) ---------- (1)

initial vertical momentum = final vertical momentum

38 * 10^-3 * 2.2= 60 * 10^-3 * v * sin(theta) ----------- (2)

on solving 1 and 2 we'll get

theta = 45 degree north of east

v = 1.9704 m/s